A football is kicked at ground level with initial velocity of 64 feet per second.
1. Y=-16t^2 + 64t
2. Y= -16(t-2)^2 + 64
3. Y=-16t(t-4)
For each of the following situations determine which form of the equation would provide with the information needed in the most efficient manner.
a. The height after 3 seconds is _____.
b. The time when the football hits the ground is _____.
C. An estimate of the time when the football is 40 feet high is ____.

Answer by Boreal(15235)

after 3 seconds, t=3 and 2 would be most efficient, although 1 is not bad.
the time it hits the ground is when -16t^2+64t=0, so when (t-4)=0 the equation is solved.
this is 16t^2-64t=0
16t(t-4)=0 or -16t(t-4)=0 and t=4 seconds so 3 is for B
estimate when football is 40 feet high would be when -16t^2+64t-40=0, and the first equation would be solved as a quadratic and be most efficient so 1 for C